OpenStudy (anonymous):

Rewrite the expression 3logx−2log(x^2+1)+4log

OpenStudy (nubeer):

hmm are u sure this is the question because i cannot see any 'x' with the last log.

OpenStudy (anonymous):

Rewrite the expression 3logx−2log(x^2+1)+4log(x-1)

OpenStudy (anonymous):

log identities: \[\log a + \log b = \log (ab)\]\[\log a - \log b = \log (a/b)\]\[a \log b = \log b^{a}\]

OpenStudy (anonymous):

Apply the 3rd identity, then the 2nd, then the 1st.

OpenStudy (anonymous):

i cant seem to find the right ans still

OpenStudy (anonymous):

What have you tried?

OpenStudy (anonymous):

all steps

OpenStudy (anonymous):

Post the expression you got after the first step you took.